### v1

# A Nonparametric Discrete Hawkes Model with a Collapsed Gaussian-Process Prior

Abstract
Hawkes process models are used in settings where past events increase the likelihood of future events occurring. Many applications record events as counts on a regular grid, yet discrete-time Hawkes models remain comparatively underused and are often constrained by fixed-form baselines and excitation kernels. In particular, there is a lack of flexible, nonparametric treatments of both the baseline and the excitation in discrete time. To this end, we propose the Gaussian Process Discrete Hawkes Process (GP-DHP), a nonparametric framework that places Gaussian process priors on both the baseline and the excitation and performs inference through a collapsed latent representation. This yields smooth, data-adaptive structure without prespecifying trends, periodicities, or decay shapes, and enables maximum a posteriori (MAP) estimation with near-linear-time $O(T\log T)$ complexity. A closed-form projection recovers interpretable baseline and excitation functions from the optimized latent trajectory. In simulations, GP-DHP recovers diverse excitation shapes and evolving baselines. In case studies on U.S. terrorism incidents and weekly Cryptosporidiosis counts, it improves test predictive log-likelihood over standard parametric discrete Hawkes baselines while capturing bursts, delays, and seasonal background variation. The results indicate that flexible discrete-time self-excitation can be achieved without sacrificing scalability or interpretability.
1 Introduction
Hawkes processes are self-exciting stochastic models in which the probability of an event occurring increases in response to past occurrences. Originally introduced by (hawkes1971, ), they have been widely adopted across disciplines such as seismology (ogata1988, ), finance (bacry2015, ), criminology (mohler2011, ), social networks (zhou2013, ; du2016, ), and epidemiology (reinhart2018, ; browning2021, ). Their appeal lies in the ability to model bursty, temporally clustered behavior via an additive decomposition of the event rate into baseline and excitation components.
Much of the literature has focused on continuous-time Hawkes processes, where the conditional intensity evolves over real-valued time and is typically expressed as the sum of a deterministic baseline and a parametric excitation kernel. Common excitation functions include exponential and power-law decay due to their simplicity and interpretability (ogata1988, ). More recent work proposes flexible alternatives using histogram-based kernels (lewis2011, ), basis expansions (zhou2013, ), neural network parameterizations (mei2017, ), and Gaussian process priors (adams2009, ; lloyd2015, ), with fully Bayesian approaches for latent network inference (linderman2014, ). In particular, Gaussian process based intensities have been developed in continuous time (e.g. (malemshinitski2021, )).
In many real-world scenarios, however, events are recorded at fixed intervals. Examples include weekly case counts in public health and daily incident logs in security contexts. Applying continuous-time Hawkes models to such data requires time discretization, which can introduce bias and hinder interpretability. Discrete Hawkes processes (DHPs) provide a principled alternative: they define event intensity directly over discrete time steps while retaining the usual self-exciting structure. Despite these advantages, most existing DHP models rely on restrictive parametric assumptions. The baseline intensity is often modeled as constant or sinusoidal, while excitation kernels are restricted to geometric or negative-binomial forms. While effective in some settings, such assumptions can limit the ability to capture long memory, nonstationarity, or changes in excitation over time.
To address these limitations, (browning2022, ) introduced a nonparametric DHP using a random histogram prior over the excitation kernel with trans-dimensional MCMC for inference. While this approach allows data-driven excitation structure, the use of a fixed intercept for the baseline may limit its ability to capture smooth or evolving background dynamics. Moreover, the sampling scheme can be costly for long processes.
In this paper, we introduce the Gaussian Process Discrete Hawkes Process (GP-DHP), a nonparametric model for discrete-time self-exciting count data. GP-DHP places independent Gaussian process priors on the baseline and excitation functions, enabling smooth, data-adaptive estimation without strong parametric constraints. Crucially, we collapse these priors to a single latent GP over the additive intensity, enabling efficient MAP inference and an interpretable decomposition into exogenous (baseline) and endogenous (excitation) components. This yields near-linear-time $\mathcal{O}(T\log T)$ complexity in practice via FFT-based multiplications and structured kernel interpolation. To our knowledge, there is currently no analogous GP-based formulation for discrete-time Hawkes that models both baseline and excitation nonparametrically; our work fills this gap.
Beyond introducing GP-DHP and the collapsed latent-GP inference described above, we also contribute:
  1. 1.
    A closed-form projection from the latent trajectory to interpretable baseline/excitation components;
  2. 2.
    Practical identifiability and stability diagnostics, including the branching-ratio statistic $\widehat{\kappa}=\sum_{d=1}^{D_{\max}}\max\{\widehat{f}(d),0\}$.
  3. 3.
    Empirical evidence on simulations and two case studies (U.S. terrorism; weekly Cryptosporidiosis) showing improved test predictive log-likelihood and faithful decompositions; and
  4. 4.
    an open-source implementation to facilitate replication and reuse.
2 Proposed Model
2.1 Discrete Hawkes Process
The discrete-time Hawkes process (DHP) is the discrete analogue of the continuous-time Hawkes process originally introduced by (hawkes1971, ). In continuous time, the intensity function defines the instantaneous rate of event occurrence and typically increases immediately following a self-exciting or clustering phenomenon. In contrast, the DHP operates over discrete time steps, $t\in\mathbb{N}$, and models the expected number of events per interval using a combination of exogenous and endogenous contributions.
The DHP is designed for count data observed at regular time intervals, such as hourly, daily, or weekly series. It captures the intuition that recent events tend to increase the probability of future events, a property known as self-excitation. This makes DHPs particularly useful for modeling clustered sequences, such as infectious disease cases, financial transactions, or crime incidents, where new occurrences can trigger follow-up events in subsequent intervals.
Let $N(t)\in\mathbb{N}$ denote the number of events observed during the interval (t1,t](t-1,t], and let the event history up to time t1t-1 be denoted $\mathcal{H}(t-1)=\{N(s):s<t\}$. The conditional intensity function, or expected event rate at time $t$, is defined as:
$\lambda(t)=\mathbb{E}[N(t)\mid\mathcal{H}(t-1)]=\mu(t)+\sum_{d=1}^{t-1}N(t-d)\,\Phi(d),$ (1)
where:
  • $\mu(t)>0$, for $t\in\mathbb{N}$, is the baseline intensity, representing spontaneous (i.e., exogenous) events not triggered by past observations. It can capture systematic background variation beyond the influence of past events. In practice, this often reflects smooth trends or seasonal cycles. For example, in epidemiology $\mu(t)$ may encode annual patterns in disease incidence, while in security contexts it may capture differences between weekdays and weekends. This separation allows the model to distinguish predictable rhythms from bursty self-excitation.
  • $\Phi(d)\in\mathbb{R}$ (non-negative in the standard case) is the excitation kernel, describing the influence that events occurring $d$ steps in the past exert on the present rate. For discrete Hawkes processes, stability requires the excitation kernel to be absolutely summable with total branching ratio $\kappa=\sum_{d=1}^{\infty}\Phi(d)<1$. This condition prevents explosive growth of events and guarantees a well-defined mean process. If signed kernels are allowed, stability is enforced by bounding the sum of the positive part, $\sum_{d\geq 1}\max\{\Phi(d),0\}<1$.
The observed counts are modeled using a Poisson distribution:
$N(t)\sim\mathrm{Poisson}(\lambda(t)).$
This likelihood is widely used in DHPs because it naturally models integer-valued events and supports the additive decomposition of the intensity function. We emphasize that this is a modeling assumption, not a consequence of discretizing the continuous-time process. That is, a Poisson likelihood is not implied by discretizing a continuous-time Hawkes process; rather, it is a standard choice in the discrete-time literature for tractability and interpretability. Our nonparametric approach refers to the structure of the intensity function rather than the observation model.
Figure 1: Discrete-time Hawkes intensity on $[0,20]$ with geometric excitation. The plot shows intensity (black step), baseline (red dashed), and event times (black dotted; multiplicities annotated at the top). Parameters: baseline $\mu=0.5$, jump size $K=0.75$, geometric decay $\beta=0.5$; events at $t=3,7,10$ with multiplicities $2,4,3$, respectively.
2.1.1 Branching Representation of the Discrete Hawkes Process
The discrete Hawkes model above is often interpreted through the Hawkes cluster representation: events arise either as immigrants from the baseline $\mu(t)$ or as offspring triggered by past events, with integer lags distributed according to $\Phi(d)$ (hawkes1971, ; hawkes1974, ). In this view, bursts form around immigrant arrivals as offspring generate further descendants. We call this the branching process interpretation.
This branching process interpretation, in some cases, facilitates practical inference and simulation. Each event at time $t$ is assigned a parent source, either an earlier event (endogenous or triggered) or a spontaneous immigrant (exogenous event). The set of all such parent–child relationships is known as the branching structure.
Let $N(t)\in\mathbb{N}$ denote the observed number of events at time $t$, and define a latent vector $\boldsymbol{n}^{(t)}=(n_{0}^{(t)},n_{1}^{(t)},\ldots,n_{t-1}^{(t)})$, where: - $n_{0}^{(t)}$ is the number of immigrant (baseline) events at time $t$, - $n_{d}^{(t)}$ is the number of offspring events at time $t$ triggered by events at time tdt-d, for $d\in\{1,\dots,t-1\}$.
These quantities satisfy the constraint:
$N(t)=\sum_{d=0}^{t-1}n_{d}^{(t)}.$
Given the branching structure $\{\boldsymbol{n}^{(t)}\}_{t=1}^{T}$, the complete-data likelihood factorizes into a product of contributions from baseline and excitation sources. The probability of an event being triggered by a parent tdt-d is proportional to the weight $\Phi(d)$, and the probability of being an immigrant is proportional to $\mu(t)$. The normalizing constant is the total expected intensity at time $t$, that is, $\lambda(t)=\mu(t)+\sum_{d=1}^{t-1}N(t-d)\Phi(d)$.
For inference, the branching assignments can be interpreted as draws from a multinomial distribution:
$(n_{0}^{(t)},n_{1}^{(t)},\ldots,n_{t-1}^{(t)})\sim\mathrm{Multinomial}\left(N(t);\frac{\mu(t)}{\lambda(t)},\frac{N(t-1)\Phi(1)}{\lambda(t)},\ldots,\frac{N(1)\Phi(t-1)}{\lambda(t)}\right).$
This formulation enables efficient simulation and inference schemes, such as Gibbs sampling, by marginalizing over the latent branching structure or sampling it explicitly.
While Gibbs sampling is effective for parametric DHPs, it can perform poorly for the model in this paper due to slow mixing under Gaussian-process priors. Thus, we develop an alternative MAP-based inference procedure that scales efficiently and avoids these issues.
2.2 Proposed Model
Standard inference in Hawkes processes typically involves estimating the baseline and excitation functions either through maximum likelihood using parametric forms, or via sampling-based Bayesian approaches. In the discrete-time setting, parametric models often assume constant or periodic baselines and simple excitation forms (e.g., exponential decay (browning2021, )), which limit flexibility. Fully Bayesian approaches with nonparametric priors exist but tend to be computationally intensive, especially when jointly estimating multiple latent functions. In contrast, our approach collapses the GP priors over baseline and excitation into a single prior over the latent intensity trajectory, enabling efficient MAP inference with scalable optimization. This not only improves interpretability but also avoids expensive sampling or marginalization steps, making the method suitable for large-scale temporal data. To this end, we propose a nonparametric extension of the DHP that models both components using Gaussian process (GP) priors. While several works have explored GP-based intensity modeling in the continuous-time Hawkes framework, see for example adams2009 ; lloyd2015 ; malemshinitski2021 , to our knowledge, there are currently no analogous GP-based formulations for discrete-time Hawkes processes. Our work addresses this gap by developing a discrete-time, nonparametric model that retains the interpretability of parametric counterparts. We first introduce the Gaussian Process.
2.2.1 Gaussian processes
A Gaussian process (GP) is a Bayesian nonparametric prior over functions. Formally, it is a collection of random variables such that any finite subset is jointly Gaussian. A GP is specified by a mean function $m(x)$ and a covariance (kernel) function $k(x,x^{\prime})$, denoted
$f(x)\sim\mathcal{GP}\big(m(x),\,k(x,x^{\prime})\big).$
For inputs $\{x_{1},\dots,x_{n}\}$, the vector of evaluations $\mathbf{f}=[f(x_{1}),\dots,f(x_{n})]^{\top}$ satisfies
$\mathbf{f}\sim\mathcal{N}(\mathbf{m},\mathbf{K}),\quad\text{with }\mathbf{m}_{i}=m(x_{i}),\;\mathbf{K}_{ij}=k(x_{i},x_{j}).$
GPs are well suited to modeling time-varying functions when the functional form is unknown but expected to exhibit structure such as smoothness, periodicity, or trends. Appropriate kernel choices encode these assumptions while keeping inference tractable; examples include the squared-exponential (RBF) kernel for smoothness, periodic kernels for seasonality, and linear kernels for trends.
In what follows, we place GP priors on both the baseline $b(t)$ and the excitation kernel $f(d)$ of the discrete Hawkes model. This allows interpretable, data-driven intensity dynamics without committing to fixed parametric forms. For a comprehensive introduction, see (rasmussen2006gaussian, ).
2.2.2 Gaussian Process Discrete Hawkes Process (GP-DHP)
We model the latent intensity $\ell$ as a discrete Hawkes process (as in Equation (1)). That is, the sum of two components: a baseline function $b(t)$ and a time-lagged excitation term parameterized by a function $f(d)$,
$\ell(t)\;=\;b(t)\;+\;\sum_{d=1}^{t-1}N(t-d)\,f(d),$ (2)
where $N(t)$ denotes the observed event count in interval (t1,t](t-1,t], and $d\in\{1,2,\dots\}$ indexes discrete lags. We encode long-lag attenuation and smoothness directly in the GP prior for $f$ (see below). To obtain a nonnegative intensity, we apply a rectifying link to the latent process:
$\lambda(t)=\max\{0,\ell(t)\},\qquad N(t)\sim\mathrm{Poisson}\big(\lambda(t)\big).$
We use this rectifier throughout. The non-differentiability at 0 can, in principle, slow gradient methods, however we found in practice that our MAP optimisation with damping and line search converges reliably.
As in any additive Hawkes decomposition, $b(t)$ and the excitation contribution $\sum_{d}N(t-d)f(d)$ are not uniquely identifiable without additional structure. If $b(t)$ were modeled by an unconstrained GP, slowly varying temporal patterns could be attributed either to the baseline or to accumulated excitation, leading to ambiguous decompositions. We resolve this by (i) imposing a seasonally structured prior on the baseline and (ii) using a nonstationary, lag-dependent prior on $f(d)$ that shrinks long-range effects. Together these priors restrict the solution space and yield stable, interpretable decompositions (see Section 4.1).
GP prior for the baseline $b(t)$.
We place a zero-mean GP prior $b\sim\mathcal{GP}(0,K_{b})$ with a kernel that captures seasonal recurrence, and along-term trend:
$K_{b}(t,t^{\prime})$ $=\underbrace{\sigma_{\mathrm{per}}^{2}\exp\!\left(-\frac{2\,\sin^{2}\!\big(\pi(t-t^{\prime})/P\big)}{\ell_{\mathrm{per}}^{2}}\right)}_{\text{periodic (seasonal)}}\;+\;\underbrace{\sigma_{\mathrm{lin}}^{2}\,tt^{\prime}}_{\text{linear (trend)}}\;$ (3)
$\hskip 20.00003pt+\;\epsilon_{b}^{2}\,\delta_{tt^{\prime}}\quad\text{(jitter for numerical stability).}$
Here $P$ is the seasonal period (e.g., P=52P{=}52 for weekly data or P=365P{=}365 for daily data), $\ell_{\mathrm{per}}$ controls seasonal smoothness, $\sigma_{\mathrm{lin}}^{2}$ the magnitude of long-term linear variation.
GP prior for the excitation kernel $f(d)$.
We place a zero-mean GP prior $f\sim\mathcal{GP}(0,K_{f})$ on discrete lags $d\in\{1,\dots,D_{\max}\}$. Rather than multiplying by an explicit exponential envelope, we encode nonstationarity through both an amplitude envelope and an input warping of the lags. Let
$a(d)\;=\;\sigma_{f}\exp\!\left(-\frac{\beta d}{2}\right),\qquad g(d)\;=\;\frac{1-e^{-\beta d}}{\beta\,\ell_{f}},$
with $\sigma_{f}>0$ an amplitude scale, $\ell_{f}>0$ a base length-scale, and $\beta\geq 0$ a parameter controlling the rate at which the effective metric and amplitude change with lag. Define the stationary squared-exponential kernel on the warped inputs by
$k_{\mathrm{RBF}}\big(g(d),g(d^{\prime})\big)\;=\;\exp\!\left(-\tfrac{1}{2}\,\big(g(d)-g(d^{\prime})\big)^{2}\right).$
Our excitation kernel is then
$K_{f}(d,d^{\prime})\;=\;a(d)\,a(d^{\prime})\;k_{\mathrm{RBF}}\big(g(d),g(d^{\prime})\big)\;+\;\epsilon_{f}^{2}\,\delta_{dd^{\prime}}.$ (4)
This construction yields smooth, short-lag correlations while simultaneously shrinking long-lag variability via $a(d)$ and compressing large lags through the warp $g(\cdot)$. As $\beta\to 0$, we recover a stationary RBF prior on the original lag scale because $a(d)\!\to\!\sigma_{f}$ and $g(d)\!\to\!d/\ell_{f}$. For $\beta>0$, the prior increasingly attenuates and decorrelates remote lags, which discourages the excitation from absorbing slow-moving trends that should be attributed to the baseline process $b(t)$.
Figure 2 illustrates how the excitation function’s GP prior behaves as the attenuation parameter $\beta$ increases. The top row shows sample draws of the excitation function $f(d)$ under different $\beta$ values, highlighting how larger $\beta$ suppresses long-lag fluctuations and concentrates mass near the origin. The bottom row displays the corresponding covariance heatmaps $K_{f}$. As $\beta$ increases, the excitation functions flatten at longer lags and exhibit smoother decay, reflecting a reduced influence of distant past events.
Discussion.
The combination of (3) and (4) provides the necessary inductive bias for identifiability: repeated seasonal structure and trend are captured by $b(t)$, whereas $f(d)$ explains short- to medium-range self-excitation with a principled decay built into the prior malemshinitski2021 (i.e., without an explicit Gaussian envelope factor $e^{-\alpha d}$ multiplied the excitation function).
(a) $f(d)$ draws, $\beta=0.00$, $\ell_{f}=10$
(b) $f(d)$ draws, $\beta=0.03$, $\ell_{f}=10$
(c) $f(d)$ draws, $\beta=0.10$, $\ell_{f}=10$
(d) $K_{f}$ covariance heatmap, $\beta=0.00$, $\ell_{f}=10$
(e) $K_{f}$ covariance heatmap, $\beta=0.03$, $\ell_{f}=10$
(f) $K_{f}$ covariance heatmap, $\beta=0.10$, $\ell_{f}=10$
Figure 2: Draws of the GP prior over the excitation function $f$: the effect of increasing $\beta$ at fixed $\ell_{f}$. Top: draws of $f(d)$; Bottom: corresponding heatmaps for the covariance matrices $K_{f}$.
3 Inference
Our inference procedure centers on the latent intensity function $l(t)$, defined in Equation (2), which we treat as the primary object of estimation. Since both the baseline $b(t)$ and the excitation kernel $f(d)$ are endowed with Gaussian process priors, the resulting $l(t)$ inherits a well-defined Gaussian process prior induced by these components (rasmussen2006gaussian, ). Rather than estimating $b(t)$ and $f(d)$ jointly, we integrate them out analytically to form a collapsed prior over $l(t)$. This yields a single GP prior over the entire latent trajectory, improving computational efficiency while preserving the underlying structure.
While discrete Hawkes models are often interpreted using a branching structure, we do not adopt this representation for inference. In our experiments, we found that standard sampling-based approaches using the branching formulation mix poorly when combined with GP priors. This motivates our use of a collapsed latent-intensity representation and direct MAP estimation, which avoids explicit branching variables and enables stable, scalable inference.
3.1 Collapsed GP Prior
[…]
Recall from Equation (2), the latent intensity at each time point is given by
$\ell(t)\;=\;b(t)\;+\;\sum_{d=1}^{t-1}N(t-d)\,f(d).$
[…]
$K(t,s)\;=\;K_{b}(t,s)\;+\;\sum_{d=1}^{t-1}\sum_{d^{\prime}=1}^{s-1}N(t-d)\,N(s-d^{\prime})\,K_{f}(d,d^{\prime}).$ (5)
Now, let $X\in\mathbb{R}^{T\times(T-1)}$ be the strictly lower–triangular “lagged-count” design matrix with
$X_{t,d}\;=\;\begin{cases}N(t-d),&1\leq d\leq t-1,\\ 0,&\text{otherwise},\end{cases}\qquad t=1,\dots,T,\ \ d=1,\dots,T-1,$
and let $K_{f}\in\mathbb{R}^{(T-1)\times(T-1)}$ be the excitation-kernel matrix with entries $K_{f}(d,d^{\prime})$ for lags $d,d^{\prime}\geq 1$. Then the collapsed prior can be written compactly as
$K\;=\;K_{b}\;+\;X\,K_{f}\,X^{\top}.$
This formulation avoids the need to represent $b(t)$ and $f(d)$ explicitly during inference. This is analogous to GP-modulated Poisson processes in continuous time malemshinitski2021 , adapted here to the discrete-time Hawkes setting with a collapsed prior. Unlike malemshinitski2021 , which imposes an explicit exponential envelope together with a stationary RBF kernel for the excitation (i.e., contributions of the form $e^{-\alpha d}$ and covariance of the GP prior of the excitation function $k_{f}(d,d^{\prime})=\sigma_{f}^{2}\exp\!\{-(d-d^{\prime})^{2}/(2\ell_{f}^{2})\}$), we remove the deterministic envelope from the mean structure and encode attenuation and smoothness inside the GP prior for $f$.
3.2 MAP Objective
To infer the latent intensity trajectory $\mathbf{l}=[l(1),\dots,l(T)]^{\top}$, we adopt a maximum a posteriori (MAP) approach. Let $\mathbf{N}=[N(1),\dots,N(T)]^{T}$ denote the observed event counts. The MAP estimate of $\mathbf{l}$, which we denote $\mathbf{l}^{*}$, maximizes the log-posterior:
$\log p(\mathbf{l}\mid\mathbf{N})\propto\log p(\mathbf{N}\mid\mathbf{l})-\frac{1}{2}\mathbf{l}^{\top}K^{-1}\mathbf{l},$
where the first term is the Poisson log-likelihood and the second is the GP prior penalty. The likelihood is given by:
$\log p(\mathbf{N}\mid\mathbf{l})=\sum_{t=1}^{T}\left(N(t)\log\lambda(t)-\lambda(t)-\log N(t)!\right),\quad\text{with}\quad\lambda(t)=\max\{0,l(t)\}.$
In optimization we drop the constant $\sum_{t}\log N(t)!$ and solve:
$\mathbf{l}^{*}=\arg\max_{\mathbf{l}}\left\{\log p(\mathbf{N}\mid\mathbf{l})-\frac{1}{2}\mathbf{l}^{\top}K^{-1}\mathbf{l}\right\}.$
3.3 Decomposition into Baseline and Excitation Components
While inference is performed on the latent intensity $\mathbf{l}^{*}$, recovering the individual components $b(t)$ and $f(d)$ is desirable for interpretability, particularly in epidemiological or social applications where disentangling baseline dynamics from self-excitation reveals mechanistic insight.
To this end, we approximate the MAP estimate of the latent intensity as a linear combination:
$\mathbf{l}^{*}\approx\mathbf{b}+\mathbf{X}\mathbf{f},$
where $\mathbf{b}\in\mathbb{R}^{T}$ is the baseline vector, $\mathbf{f}\in\mathbb{R}^{T-1}$ is the excitation kernel evaluated at discrete lags, and $\mathbf{X}\in\mathbb{R}^{T\times(T-1)}$ is a fixed design matrix. Each row of $\mathbf{X}$ encodes the weighted contributions of past events, defined as:
$\mathbf{X}[t,d]=\begin{cases}N(t-d),&\text{if }d<t,\\ 0,&\text{otherwise}.\end{cases}$
This decomposition mirrors the additive structure of Equation (2), rewritten in matrix form. It provides a natural way to project the latent intensity $\mathbf{l}^{*}$ back onto interpretable baseline and excitation components.
Ideally, one would recover $\mathbf{b}$ and $\mathbf{f}$ by solving the constrained optimization problem:
$\min_{\mathbf{b},\,\mathbf{f}}\left\{\frac{1}{2}\mathbf{b}^{\top}K_{b}^{-1}\mathbf{b}+\frac{1}{2}\mathbf{f}^{\top}K_{f}^{-1}\mathbf{f}\right\}\quad\text{subject to }\mathbf{l}^{*}=\mathbf{b}+\mathbf{X}\mathbf{f}.$
This corresponds to the maximum a posteriori estimate under the constraint that $\mathbf{l}^{*}$ is exactly decomposed into baseline and excitation terms. We next show that there exists a unique closed-form solution to this constrained optimisation problem.
Proposition 3.1 (Hard-constraint decomposition: existence, uniqueness, and closed form solution).
Let $K_{b}\in\mathbb{R}^{T\times T}$ and $K_{f}\in\mathbb{R}^{(T-1)\times(T-1)}$ be symmetric positive definite, let $X\in\mathbb{R}^{T\times(T-1)}$, and let $\ell^{*}\in\mathbb{R}^{T}$. Consider
$\min_{b,f}\;\frac{1}{2}\,b^{\top}K_{b}^{-1}b+\frac{1}{2}\,f^{\top}K_{f}^{-1}f\quad\text{subject to}\quad\ell^{*}=b+Xf.$ (6)
Define $K:=K_{b}+XK_{f}X^{\top}\in\mathbb{R}^{T\times T}$. Then:
  1. 1.
    The feasible set is nonempty and the objective is strictly convex on it. Hence there is a unique minimizer.
  2. 2.
[…]
    $\widehat{b}\;=\;K_{b}K^{-1}\ell^{*},\qquad\widehat{f}\;=\;K_{f}X^{\top}K^{-1}\ell^{*}.$
  3. 3.
    The minimum value equals $\tfrac{1}{2}\,\ell^{*\top}K^{-1}\ell^{*}$.
Proof.
Feasibility and strict convexity. Feasibility holds since $(b,f)=(\ell^{*},0)$ satisfies $\ell^{*}=b+Xf$. The objective $J(b,f)=\tfrac{1}{2}b^{\top}K_{b}^{-1}b+\tfrac{1}{2}f^{\top}K_{f}^{-1}f$ is strictly convex because $K_{b}^{-1}\succ 0$ and $K_{f}^{-1}\succ 0$. The feasible set is an affine subspace. Therefore a unique minimizer exists.
KKT conditions and solution. Form the Lagrangian
$\mathcal{L}(b,f,\lambda)=\tfrac{1}{2}b^{\top}K_{b}^{-1}b+\tfrac{1}{2}f^{\top}K_{f}^{-1}f+\lambda^{\top}(\ell^{*}-b-Xf),$
with multiplier $\lambda\in\mathbb{R}^{T}$. Stationarity gives
$\nabla_{b}\mathcal{L}=K_{b}^{-1}b-\lambda=0\;\Rightarrow\;b=K_{b}\lambda,\qquad\nabla_{f}\mathcal{L}=K_{f}^{-1}f-X^{\top}\lambda=0\;\Rightarrow\;f=K_{f}X^{\top}\lambda.$
Primal feasibility enforces
$\ell^{*}-b-Xf=\ell^{*}-K_{b}\lambda-XK_{f}X^{\top}\lambda=0\;\;\Rightarrow\;\;K\lambda=\ell^{*},$
where $K:=K_{b}+XK_{f}X^{\top}$. For any nonzero $v\in\mathbb{R}^{T}$,
$v^{\top}Kv=v^{\top}K_{b}v+(X^{\top}v)^{\top}K_{f}(X^{\top}v)>0,$
since $K_{b}\succ 0$ and $K_{f}\succ 0$. Hence $K\succ 0$ and is invertible. Thus $\lambda=K^{-1}\ell^{*}$, and substituting back yields the claimed closed forms
$\widehat{b}=K_{b}K^{-1}\ell^{*},\qquad\widehat{f}=K_{f}X^{\top}K^{-1}\ell^{*}.$
These satisfy the KKT system and the constraint, hence are optimal. Uniqueness follows from strict convexity.
Minimum value. Using $\widehat{b}=K_{b}\lambda$ and $\widehat{f}=K_{f}X^{\top}\lambda$ with $\lambda=K^{-1}\ell^{*}$,
$J(\widehat{b},\widehat{f})=\tfrac{1}{2}\lambda^{\top}K_{b}\lambda+\tfrac{1}{2}\lambda^{\top}XK_{f}X^{\top}\lambda=\tfrac{1}{2}\lambda^{\top}K\lambda=\tfrac{1}{2}\,\ell^{*\top}K^{-1}\ell^{*}.$
This completes the proof. ∎
[…]
3.4 Computational Complexity and Efficiency
A key advantage of our framework lies in its ability to scale to long sequences while maintaining tractable computational cost. Classical implementations of discrete Hawkes processes, as defined in Equation (1), often incur $\mathcal{O}(T^{2})$ complexity. This is because evaluating the excitation term
$\sum_{d=1}^{t-1}N(t-d)\,\Phi(d)$
at each time $t$ requires summing over the full event history, and this must be repeated for each time step $t=1,\dots,T$. While truncation strategies can reduce this burden in parametric settings, such techniques are less effective in nonparametric models due to the bias introduced by ignoring long-range dependencies.
In contrast, our GP-DHP formulation allows for scalable inference, with efficient evaluation of the log-likelihood and gradients during optimization. This is primarily due to two key design features:
  • Collapsed latent GP formulation: Rather than placing priors on the baseline and excitation functions separately and sampling or optimizing them directly, we collapse these into a single Gaussian Process prior over the latent additive intensity $\ell(t)$, as described in Equation (5). This enables MAP estimation over the latent function directly, avoiding repeated explicit recomputation of excitation sums at each time step.
  • Efficient log-likelihood evaluation: With a Poisson observation model and an additive latent intensity, the log-likelihood decomposes into a sum over $t=1,\dots,T$, where each term depends only on $\ell(t)$. Temporal dependencies are encoded in the GP prior, so the likelihood evaluation itself contains no nested history summations.
An additional computational benefit arises from the structure of the joint GP kernel used during inference. The excitation contribution to the covariance can be written as
$K_{\text{exc}}(t,s)$ $=\sum_{d=1}^{t-1}\sum_{d^{\prime}=1}^{s-1}N(t-d)\,N(s-d^{\prime})\,K_{f}(d,d^{\prime})$
$=X\,K_{f}\,X^{\top},$
where $X\in\mathbb{R}^{T\times(T-1)}$ is the strictly lower–triangular “lagged-count” design matrix with $X_{t,d}=N(t-d)$ for $d<t$ and 0 otherwise, and $K_{f}$ is the excitation-kernel matrix with entries $K_{f}(d,d^{\prime})$. Under our model, $K_{f}$ admits the factorization
$K_{f}\;=\;A\,K_{\text{stat}}\,A,\qquad A:=\mathrm{diag}\!\big(a(1),\dots,a(T\!-\!1)\big),\quad a(d)=\sigma_{f}e^{-\beta d/2},$
with a stationary RBF kernel applied to the warped lags $u(d)=g(d)=(1-e^{-\beta d})/(\beta\,\ell_{f})$:
$\big[K_{\text{stat}}\big]_{d,d^{\prime}}\;=\;\exp\!\Big(-\tfrac{1}{2}\,[u(d)-u(d^{\prime})]^{2}\Big).$
Hence
$K_{\text{exc}}\;=\;(XA)\,K_{\text{stat}}\,(XA)^{\top}.$
Fast matrix-vector multiplications (MVMs).
MAP inference with a GP prior requires solving linear systems of the form $K\,\boldsymbol{v}$, so efficiency hinges on fast MVMs:
  1. 1.
    Multiplication by $X$ and $X^{\top}$ as convolutions. For any vector $\boldsymbol{y}$, the products $X^{\top}\boldsymbol{y}$ and $X\boldsymbol{w}$ are (cross-)correlations between $\boldsymbol{y}$ (or $\boldsymbol{w}$) and the event-count sequence $N(\cdot)$. These can be computed in $\mathcal{O}(T\log T)$ time via FFT-based linear convolution (after standard zero-padding), and the FFT of $N(\cdot)$ can be reused across iterations.
  2. 2.
    Multiplication by $A$ is diagonal. Left/right multiplication by $A$ costs $\mathcal{O}(T)$.
  3. 3.
    Multiplication by $K_{\text{stat}}$. Although $K_{\text{stat}}$ is stationary in the warped input $u(d)$, it is not Toeplitz with respect to the integer lag index $d$. We therefore use a structured-kernel–interpolation (inducing-grid) approximation: choose $M\ll T$ inducing points on a uniform grid in the $u$-domain, form a sparse interpolation matrix $W\in\mathbb{R}^{(T-1)\times M}$, and approximate
    $K_{\text{stat}}\;\approx\;W\,K_{U}\,W^{\top},$
    where $K_{U}$ is the stationary RBF kernel on the uniform grid. Because the grid is uniform, $K_{U}$ is (block-)Toeplitz/circulant and supports $\mathcal{O}(M\log M)$ MVMs via FFT (see xia2007toeplitz ). Since $W$ is very sparse, $W\boldsymbol{v}$ and $W^{\top}\boldsymbol{v}$ cost $\mathcal{O}(T)$.
Combining these steps yields the following per–CG-iteration complexity for the excitation block:
$\boldsymbol{y}\;\mapsto\;K_{\text{exc}}\boldsymbol{y}\;=\;XA\,\underbrace{(WK_{U}W^{\top})}_{\approx K_{\text{stat}}}\!AX^{\top}\boldsymbol{y}\quad\Rightarrow\quad\mathcal{O}\big(T\log T\big)\;+\;\mathcal{O}\big(M\log M\big)\;+\;\mathcal{O}(T).$
The baseline block $K_{b}$ is the sum of a periodic kernel (amenable to FFT/circulant embedding for $\mathcal{O}(T\log T)$ MVMs) and rank-one terms (linear and constant), so $K_{b}\boldsymbol{v}$ also costs $\mathcal{O}(T\log T)$. In practice, the overall per-iteration cost is dominated by FFTs on vectors of length $\mathcal{O}(T)$ (and $\mathcal{O}(M)$ for the inducing grid), and memory is linear in $T$ plus the storage for $W$ (proportional to its number of nonzeros).
Uncertainty via Laplace and projection.
[…]
$H\;=\;K^{-1}\;+\;D,\qquad D\;=\;\mathrm{diag}\!\Big(\tfrac{N(t)}{\lambda(t)^{2}}\,\mathbf{1}\{\ell(t)>0\}\Big)_{t=1}^{T},$
[…]
$\mathrm{Cov}\!\begin{pmatrix}\widehat{b}\\ \widehat{f}\end{pmatrix}\;\approx\;\begin{pmatrix}P_{b}\\ P_{f}\end{pmatrix}\Sigma_{\ell}\begin{pmatrix}P_{b}\\ P_{f}\end{pmatrix}^{\!\top}$,
yielding pointwise bands from the corresponding marginals; in practice we report posterior means and 95% intervals for $b$ and $f$ computed from these marginals.
Identifiability and regularization.
Without additional structure the model is not identifiable: both $b(t)$ and the excitation term $\sum_{d=1}^{t-1}N(t-d)\,f(d)$ enter additively, and the effective number of latent degrees of freedom grows with $T$. In our formulation, identifiability is improved by: (i) the periodic component of $K_{b}$, which captures recurrent seasonal patterns; (ii) the linear and constant components, which absorb long-term level and trend; and (iii) the prior nonstationary structure $K_{f}=A\,K_{\text{stat}}\,A$, where the amplitude envelope $a(d)=\sigma_{f}e^{-\beta d/2}$ down-weights distant lags and the warp $g(d)=(1-e^{-\beta d})/(\beta\,\ell_{f})$ compresses the effective metric at larger lags. We do not place any explicit exponential envelope in the mean of the excitation; attenuation is induced inside the covariance through $a(\cdot)$ and $g(\cdot)$. This reduces the tendency of $f(\cdot)$ to absorb slow trends that should be attributed to $b(t)$, and encourages stable decompositions.
Stability & diagnostics.
For nonnegative excitation, a sufficient stability condition is the usual branching ratio $\kappa=\sum_{d\geq 1}f(d)<1$; for signed kernels a conservative check is $\sum_{d\geq 1}\max\{f(d),0\}<1$. Our nonstationary prior on $f$ (amplitude envelope $a(d)$ and warp $g(d)$) down-weights distant lags and thereby encourages a small branching ratio in practice. A simple post-fit diagnostic is
$\widehat{\kappa}\;=\;\sum_{d=1}^{D_{\max}}\max\{\widehat{f}(d),0\},$
computed from the projected excitation $\widehat{f}$ (Prop. 3.1).
4 Experiments
4.1 Synthetic Data
We evaluate the ability of GP-DHP to recover latent self-exciting dynamics by simulating from known discrete Hawkes models and assessing whether the model can reconstruct both the baseline and excitation components from observed count data. Neither the baseline nor excitation functions are assumed known during inference.
Model Specification
Each simulated time series is generated from a discrete Hawkes process of length $T=6000$. The event intensity is composed of a baseline function $b(t)$ and an excitation kernel $f(d)$, as described in Section 2. The baseline is modeled as a Gaussian process with a sum of periodic, linear, and constant kernels:
$K_{b}(t,t^{\prime})\;=\;\sigma_{b}^{2}\exp\!\left(-\frac{2\sin^{2}\!\big(\pi(t-t^{\prime})/P\big)}{\ell_{b}^{2}}\right)\;+\;\sigma_{b,c}^{2}\,tt^{\prime}\;,$
where $P=52$ encodes annual periodicity (e.g., for weekly data), $\ell_{b}$ controls the smoothness of the seasonal component, and $\sigma_{b,c}^{2}$ captures the magnitude of linear trend.
The excitation kernel $f(d)$ is assigned a zero-mean Gaussian Process prior that is nonstationary in lag, built from an amplitude envelope and a warped RBF:
$a(d)\;=\;\sigma_{f}\exp\!\Big(-\tfrac{\beta d}{2}\Big),\qquad g(d)\;=\;\frac{1-e^{-\beta d}}{\beta\,\ell_{f}},$
$K_{f}(d,d^{\prime})\;=\;a(d)\,a(d^{\prime})\;\exp\!\left(-\tfrac{1}{2}\,\big(g(d)-g(d^{\prime})\big)^{2}\right).$
This construction smoothly correlates nearby lags while attenuating long-range effects through $a(d)$ and compressing large lags via $g(d)$. Both GP priors are centered at zero. Inference proceeds over the latent additive intensity, followed by the closed-form projection to $b$ and $f$ (see Section 3).
Hyperparameter Selection via Cross-Validation
We select all kernel hyperparameters and the lag-attenuation parameter $\beta$ via grid-based cross-validation. The first 4,000 time steps are used as a training set, and the remaining 2,000 as a validation set. Hyperparameters are chosen to maximize the Poisson log-likelihood on the validation set.
The following grid is used during cross-validation procedures in both synthetic and real data experiments:
Hyperparameter Search Range
Lag attenuation $\beta$ {0.1, 0.2, 0.3, 0.4}
Baseline variance $\sigma_{b}$ {0.0001, 0.01, 1.0}
Baseline periodic lengthscale $\ell_{b}$ {1, 5, 100}
Linear trend variance $\sigma_{b,c}$ {0, $10^{-2}$, $10^{-4}$}
Excitation variance $\sigma_{f}$ {0.5, 1, 2}
Excitation lengthscale $\ell_{f}$ {5, 10, 20, 30}
Table 1: Hyperparameter ranges explored via grid search during cross-validation procedures in both synthetic and real data experiments.
4.1.1 Simulation Design - Excitation Function Recovery
To assess the flexibility of GP-DHP, we simulate twelve synthetic datasets from discrete Hawkes models with fixed baseline structure but varying excitation dynamics. In each case, the goal is to evaluate whether the model can accurately recover the latent components from count observations.
Baseline Structure
Across all simulations, we use a baseline function with a combination of linear and seasonal terms:
$\mu(t)=a+bt+c\sin\left(\frac{2\pi t}{P}\right)+d\cos\left(\frac{2\pi t}{P}\right),$
where $a,b,c,d,P$ are fixed constants. This form mimics typical nonstationary temporal behavior observed in epidemiological and social event data. The baseline is fitted jointly with the excitation kernel in each experiment.
Excitation Kernel Families
Each dataset varies only in the excitation kernel $f(d)$, which is drawn from one of four parametric families. In total, twelve configurations are used—three parameterizations per family:
  • Negative Binomial (NB): A heavy-tailed, overdispersed kernel with long memory:
    $f(d)=\alpha\cdot\binom{d+r-1}{d}(1-p)^{d}p^{r},$
    where $r>0$ and $p\in(0,1)$. The parameter $r$ controls peakiness, while $p$ governs tail decay.
  • Geometric: A memoryless excitation kernel (the r=1r{=}1 special case of the negative binomial):
    $f(d)=\alpha\cdot p(1-p)^{d-1},\quad d\geq 1,\ \ p\in(0,1).$
    Smaller $p$ induces longer-range dependence.
  • Power Law: A polynomially decaying kernel with tunable amplitude, near-lag width, and tail exponent:
    $f(d)=\alpha\,(\gamma+d)^{-\beta},$
    where $\alpha>0$ sets the overall scale, $\gamma\geq 0$ controls early-lag width/onset, and $\beta>1$ determines tail heaviness (larger $\beta$ implies faster decay).
  • Bimodal Gaussian Mixture: A continuous mixture of two Gaussian modes:
    $f(d)=\alpha\cdot\left[\frac{1}{2\sqrt{2\pi}\sigma}\exp\!\left(-\frac{(d-\mu_{1})^{2}}{2\sigma^{2}}\right)+\frac{1}{2\sqrt{2\pi}\sigma}\exp\!\left(-\frac{(d-\mu_{2})^{2}}{2\sigma^{2}}\right)\right],$
    with $\sigma>0$ and locations $\mu_{1},\mu_{2}$ specifying the modes. This form captures multimodal excitation with symmetric weight and width.
In each case, a time series of length $T=6,000$ is simulated, the model is fit (with the first 4,000 entries of the sequence used for training and the next 2,000 used for validation), and the projected posterior mean of the excitation kernel $\hat{f}(d)$ is compared with the ground truth. As shown in Figure 3, GP-DHP accurately recovers a broad range of excitation structures despite the presence of a shared nonstationary baseline $b(t)$.
Figure 3 presents the fitted excitation kernels for all twelve simulations, grouped by kernel family. The model consistently isolates excitation dynamics across these distinct functional forms.
(a) Negative Binomial, $(\alpha,p,r)=(0.6,0.6,2)$
(b) Negative Binomial, $(\alpha,p,r)=(0.6,0.6,4)$
(c) Negative Binomial, $(\alpha,p,r)=(0.6,0.6,6)$
(d) Geometric, $(\alpha,p)=(0.8,0.3)$
(e) Geometric, $(\alpha,p)=(0.8,0.6)$
(f) Geometric, $(\alpha,p)=(0.8,0.9)$
(g) Power, $(\alpha,\gamma,\beta)=(20,2,4)$
(h) Power, $(\alpha,\gamma,\beta)=(100,4,4)$
(i) Power, $(\alpha,\gamma,\beta)=(600,8,4)$
(j) Bimodal Gaussian,
$(\alpha,\mu_{1},\mu_{2},\sigma)=(0.8,1,6,1)$
(k) Bimodal Gaussian,
$(\alpha,\mu_{1},\mu_{2},\sigma)=(0.8,1,8,1)$
(l) Bimodal Gaussian,
$(\alpha,\mu_{1},\mu_{2},\sigma)=(0.8,1,10,1)$
Figure 3: Fitted excitation functions $\hat{f}(d)$ for twelve synthetic scenarios, grouped by kernel family. Rows correspond to: (1) Negative Binomial with increasing shape $r$ at fixed $\alpha$ and $p$; (2) Geometric with varying decay parameter $p$ at fixed $\alpha$; (3) Power law $f(d)=\alpha(\gamma+d)^{-\beta}$ with fixed $\beta=4$ and increasing width via $\gamma$ (with corresponding changes in $\alpha$); and (4) Bimodal Gaussian mixtures with fixed $\mu_{1}$ and $\sigma$ and increasing separation via $\mu_{2}$. All datasets share the same baseline $b(t)$.
4.1.2 Simulation Design - Baseline Function Recovery
Figure 4: Recovery of baseline and excitation components across three distinct baseline settings. Top panel: Comparison of true (solid blue) and estimated (dashed red) baseline functions for each experiment. The functional forms correspond to constant, linear, and periodic baselines, respectively. Bottom panel: Estimated excitation kernels (colored dashed lines) overlaid on the true excitation kernel (black solid). Despite varying baselines, the recovered excitation functions are consistent, confirming decomposition stability.
To assess the identifiability and robustness of our decomposition procedure, we conduct a controlled experiment in which the excitation kernel is held fixed (but is still estimated alongside the baseline function) while the baseline function varies. Specifically, we simulate three synthetic sequences of length $T=6000$ (first 4,000 for training; next 2,000 for validation), each using the same latent excitation function $f(d)$ drawn from a negative binomial kernel:
$f(d)=\alpha\cdot\binom{d+r-1}{d}(1-p)^{d}p^{r}.$
Only the baseline structure differs across the three cases, corresponding to three qualitatively distinct functional forms:
  1. 1.
    Constant
  2. 2.
    Linearly increasing
  3. 3.
    Linearly increasing + periodic.
After fitting GP-DHP to each dataset, we apply the post hoc decomposition described in Section 3 to recover the estimated baseline and excitation components. Figure 4 presents the results. In all three cases, the model accurately recovers the shape of the true baseline, and the estimated excitation functions remain stable across experiments. This demonstrates that GP-DHP is capable of effectively disentangling spontaneous activity from triggered self-excitation, even when baseline dynamics differ substantially.
As discussed in Section 3.4 we quantify uncertainty in the decomposed baseline and excitation functions by using a Laplace approximation around their MAP estimates. Specifically, we approximate the local posterior distribution of $(b,f)$ by a Gaussian centered at the MAP solution, draw samples from this approximation, and propagate them through the decomposition. Pointwise $95\%$ intervals are then obtained from the empirical quantiles across these draws, and figures display the posterior mean (dashed) together with the corresponding credible bands.
4.2 Real Data
We evaluate GP-DHP and four benchmark parametric discrete-time Hawkes models on two real-world count datasets. For the benchmarks, we fix the excitation function to a negative-binomial form (a commonly used parametric form of the excitation function (white2012terrorism, )), whereas GP-DHP models the excitation nonparametrically via a Gaussian process prior. All benchmark models share the same negative-binomial excitation kernel:
$\Phi(d)=\binom{d+r-1}{d}(1-p)^{d}p^{r},$
where $d\in\mathbb{N}$, $r>0$ is a dispersion parameter, and $p\in(0,1)$ controls temporal decay. This family encompasses both heavy-tailed and short-memory excitation. When $r=1$, the kernel reduces to the geometric distribution:
$\Phi(d)=(1-p)^{d}p,$
which has been used in recent infectious disease applications, including modeling COVID-19 mortality browning2021 . The more flexible negative binomial kernel has also been used in terrorism modeling white2012terrorism .
Each model is evaluated using a one-step-ahead predictive log-likelihood (pLL) on a held-out test set. For a time series of length $T$, with test indices $\mathcal{T}_{\mathrm{test}}\subset\{1,\dots,T\}$, we compute:
$\mathrm{pLL}=\sum_{t\in\mathcal{T}_{\mathrm{test}}}\log p(N(t)\mid\mathcal{H}(t-1)),$
where $\mathcal{H}(t-1)=\{N(s):s<t\}$ is the event history up to time t1t-1, and $p(N(t)\mid\mathcal{H}(t-1))$ is the predictive distribution under the fitted model.
Hyperparameters for GP-DHP are selected via cross-validation using the same grid as in Table 1. The benchmark models differ only in their specification of the baseline intensity function $\mu(t)$:
  • Discrete DHP: $\mu(t)=\gamma_{0}$
  • Linear DHP: $\mu(t)=\gamma_{0}+\gamma_{1}t$
  • Sinusoidal DHP: $\mu(t)=\gamma_{0}+\gamma_{1}\sin(2\pi t/P)$, with $P=52$ for Cryptosporidiosis and $P=365$ for U.S. Terrorism events.
  • Linear + Sinusoidal DHP: $\mu(t)=\gamma_{0}+\gamma_{1}t+\gamma_{2}\sin(2\pi t/P)$, with $P=52$ for Cryptosporidiosis and $P=365$ for U.S. Terrorism events.
4.2.1 U.S. Terrorism
We analyze a 21-year time series of daily terrorist incidents within the United States, drawn from the Global Terrorism Database (GTD) gtd . The data spans from 1972 onward. All events are aggregated to daily resolution and partitioned into training (1972–1981), validation (1982–1987), and test (1988–1992) intervals.
Terrorist activity over this period is characterized by long periods of inactivity interspersed with sudden bursts of incidents, often linked to domestic protest groups or separatist campaigns. This sparse and irregular structure poses a challenge for traditional autoregressive models, which lack the ability to explicitly model causal excitation between events.
Self-exciting models offer a principled alternative. Prior work by (white2012terrorism, ) applied a discrete Hawkes process with a negative binomial excitation kernel to terrorism data in Southeast Asia, demonstrating how such models can recover meaningful structure in burst-prone time series and yield interpretable quantities such as volatility and resilience.
In our application, GP-DHP achieves the highest predictive log-likelihood on the test set, outperforming all parametric baselines (Table 2). Notably, the linear DHP performs worse than the constant model, suggesting no benefit from including a long-term trend. The sinusoidal baseline yields a negligible improvement. In contrast, GP-DHP captures the underlying dynamics without imposing rigid structure, adapting flexibly to both long stretches of inactivity and short-term bursts. On top of this, the branching ratio for the excitation function was $\sum_{d\geq 1}\max\{\Phi(d),0\}=0.24$. Since the branching ratio is less than one, we get a stable process.
Model pLL
Discrete DHP -607.7
Linear DHP -631.0
Sinusoidal DHP -607.6
Linear + Sinusoidal DHP -628.1
GP-DHP -573.2
Table 2: Predictive log-likelihood (pLL) on U.S. terrorism data (test set).
(a) Estimated baseline and excitation functions.
(b) One-year view of the seasonal baseline component (Jan–Dec).
Figure 5: Estimated baseline and excitation functions for the GP-DHP fit to U.S. terrorism data, including a close-up view of the seasonality over one year (daily aggregation).
4.2.2 Cryptosporidiosis
We analyze a 365-week series of weekly Cryptosporidiosis case counts, available via the tscount package in R tscount . Cryptosporidiosis is a gastrointestinal disease caused by the protozoan parasite Cryptosporidium, which is typically spread through ingestion of contaminated water. While low-level background transmission persists in many regions, the disease is most notable for its sporadic and localized outbreaks. One of the most severe documented events occurred in Milwaukee in 1993, where over 400,000 residents were affected due to a contaminated municipal water supply hunter2003epidemiology ; yoder2010cryptosporidiosis . These outbreak dynamics are characterized by sharp, short-term increases in reported cases, followed by a rapid return to low endemic levels.
Such temporal patterns are challenging for classical autoregressive count models, such as Poisson or INGARCH-type models fokianos2009poisson ; fokianos2011log , which assume dependence based on past counts but do not directly model the self-exciting nature of outbreaks. These models often struggle to distinguish between sustained background intensity and transient, event-driven clustering.
In contrast, GP-DHP explicitly decomposes the latent event rate into two components: a baseline function and a self-exciting excitation kernel. This separation provides epidemiologically meaningful insights: the baseline reflects ongoing exposure or environmental risk, while the excitation term captures outbreak-driven propagation.
Recent work by (browning2021, ) demonstrated the use of discrete Hawkes models with geometric excitation kernels for modeling daily COVID-19 mortality in multiple countries. The geometric kernel is a special case of the negative binomial excitation used here (specifically, when $r=1$), and has proven effective for modeling short-memory self-excitation in epidemic settings. By generalizing to a full negative binomial kernel, GP-DHP can flexibly adapt to varying outbreak shapes and durations. On top of this, the branching ratio for the excitation function was $\sum_{d\geq 1}\max\{\Phi(d),0\}=0.86$. Since the branching ratio is less than one, we get a stable process.
On this dataset, GP-DHP achieves a modest improvement in predictive performance, see Table 3, compared to the larger gains on terrorism (Table 2) and produces interpretable decompositions of the observed dynamics (Figure 6). This method not only enhances predictive accuracy but also provides a principled way to distinguish irregular spikes from sustained background trends, which is essential for outbreak monitoring and forecasting in public health surveillance.
Model pLL
Discrete DHP -287.1
Linear DHP -290.1
Sinusoidal DHP -287.1
Linear + Sinusoidal DHP -349.8
GP-DHP -285.1
Table 3: Predictive log-likelihood (pLL) on Cryptosporidiosis data (test set).
Figure 6: Estimated baseline and excitation functions for the GP-DHP fit to Cryptosporidiosis data (weekly aggregation
GP-DHP attains the highest pLL, narrowly outperforming all parametric baselines. While the improvements are modest, this reflects the benefit of data-driven structure when temporal dynamics are weak or irregular.
5 Conclusion
We introduced GP-DHP, a scalable nonparametric model for discrete-time Hawkes processes that learns both baseline and excitation directly from binned count data. By collapsing the GP priors into a single latent process, GP-DHP admits efficient MAP optimization with $O(T\log T)$ cost and then uses a closed-form projection to recover interpretable components. Across controlled simulations and two applied datasets, GP-DHP consistently matches or exceeds the predictive performance of parametric baselines while yielding diagnostics and summaries that practitioners expect, including excitation strength, effective memory, and seasonal background behavior.
Practically, the model is attractive when baseline structure is unknown or nonstationary and when excitation is unlikely to follow a fixed parametric shape. The collapsed formulation makes the method usable at the time scales typical of surveillance and monitoring data, and the decomposition clarifies exogenous versus endogenous drivers for domain interpretation.
Limitations include reliance on MAP rather than full Bayesian inference and the current focus on univariate series. Future work includes multivariate and spatial extensions for interacting series, and data-driven kernel selection to adapt automatically to domain-specific seasonality and memory.

### v2

@@
 Hawkes processes are self-exciting stochastic models in which the probability of an event occurring increases in response to past occurrences.
-Originally introduced by (hawkes1971, ), they have been widely adopted across disciplines such as seismology (ogata1988, ), finance (bacry2015, ), criminology (mohler2011, ), social networks (zhou2013, ; du2016, ), and epidemiology (reinhart2018, ; browning2021, ).
+Originally introduced by (Hawkes, 1971), they have been widely adopted across disciplines such as seismology (Ogata, 1988), finance (Bacry et al., 2015), criminology (Mohler et al., 2011), social networks (Zhou et al., 2013; Du et al., 2016), and epidemiology (Reinhart, 2018; Browning et al., 2021).
 Their appeal lies in the ability to model bursty, temporally clustered behavior via an additive decomposition of the event rate into baseline and excitation components.
 Much of the literature has focused on continuous-time Hawkes processes, where the conditional intensity evolves over real-valued time and is typically expressed as the sum of a deterministic baseline and a parametric excitation kernel.
-Common excitation functions include exponential and power-law decay due to their simplicity and interpretability (ogata1988, ).
-More recent work proposes flexible alternatives using histogram-based kernels (lewis2011, ), basis expansions (zhou2013, ), neural network parameterizations (mei2017, ), and Gaussian process priors (adams2009, ; lloyd2015, ), with fully Bayesian approaches for latent network inference (linderman2014, ).
+Common excitation functions include exponential and power-law decay due to their simplicity and interpretability (Ogata, 1988).
+More recent work proposes flexible alternatives using histogram-based kernels (Lewis and Mohler, 2011), basis expansions (Zhou et al., 2013), neural network parameterizations (Mei and Eisner, 2017), and Gaussian process priors (Adams et al., 2009; Lloyd et al., 2015), with fully Bayesian approaches for latent network inference (Linderman and Adams, 2014).
 In particular, Gaussian process based intensities have been developed in continuous time (e.g.
-(malemshinitski2021, )).
+(Malem-Shinitski et al., 2022)).
 In many real-world scenarios, however, events are recorded at fixed intervals.
@@
 While effective in some settings, such assumptions can limit the ability to capture long memory, nonstationarity, or changes in excitation over time.
-To address these limitations, (browning2022, ) introduced a nonparametric DHP using a random histogram prior over the excitation kernel with trans-dimensional MCMC for inference.
+To address these limitations, (Browning et al., 2022) introduced a nonparametric DHP using a random histogram prior over the excitation kernel with trans-dimensional MCMC for inference.
 While this approach allows data-driven excitation structure, the use of a fixed intercept for the baseline may limit its ability to capture smooth or evolving background dynamics.
@@
 2.1 Discrete Hawkes Process
-The discrete-time Hawkes process (DHP) is the discrete analogue of the continuous-time Hawkes process originally introduced by (hawkes1971, ).
+The discrete-time Hawkes process (DHP) is the discrete analogue of the continuous-time Hawkes process originally introduced by (Hawkes, 1971).
 In continuous time, the intensity function defines the instantaneous rate of event occurrence and typically increases immediately following a self-exciting or clustering phenomenon.
@@
 2.1.1 Branching Representation of the Discrete Hawkes Process
-The discrete Hawkes model above is often interpreted through the Hawkes cluster representation: events arise either as immigrants from the baseline $\mu(t)$ or as offspring triggered by past events, with integer lags distributed according to $\Phi(d)$ (hawkes1971, ; hawkes1974, ).
+The discrete Hawkes model above is often interpreted through the Hawkes cluster representation: events arise either as immigrants from the baseline $\mu(t)$ or as offspring triggered by past events, with integer lags distributed according to $\Phi(d)$ (Hawkes, 1971; Hawkes and Oakes, 1974).
 In this view, bursts form around immigrant arrivals as offspring generate further descendants.
@@
 Standard inference in Hawkes processes typically involves estimating the baseline and excitation functions either through maximum likelihood using parametric forms, or via sampling-based Bayesian approaches.
-In the discrete-time setting, parametric models often assume constant or periodic baselines and simple excitation forms (e.g., exponential decay (browning2021, )), which limit flexibility.
+In the discrete-time setting, parametric models often assume constant or periodic baselines and simple excitation forms (e.g., exponential decay (Browning et al., 2021)), which limit flexibility.
 Fully Bayesian approaches with nonparametric priors exist but tend to be computationally intensive, especially when jointly estimating multiple latent functions.
@@
 To this end, we propose a nonparametric extension of the DHP that models both components using Gaussian process (GP) priors.
-While several works have explored GP-based intensity modeling in the continuous-time Hawkes framework, see for example adams2009 ; lloyd2015 ; malemshinitski2021 , to our knowledge, there are currently no analogous GP-based formulations for discrete-time Hawkes processes.
+While several works have explored GP-based intensity modeling in the continuous-time Hawkes framework, see for example Adams et al.
+(2009); Lloyd et al.
+(2015); Malem-Shinitski et al.
+(2022), to our knowledge, there are currently no analogous GP-based formulations for discrete-time Hawkes processes.
 Our work addresses this gap by developing a discrete-time, nonparametric model that retains the interpretability of parametric counterparts.
@@
 This allows interpretable, data-driven intensity dynamics without committing to fixed parametric forms.
-For a comprehensive introduction, see (rasmussen2006gaussian, ).
+For a comprehensive introduction, see (Rasmussen and Williams, 2006).
 2.2.2 Gaussian Process Discrete Hawkes Process (GP-DHP)
@@
 Discussion.
-The combination of (3) and (4) provides the necessary inductive bias for identifiability: repeated seasonal structure and trend are captured by $b(t)$, whereas $f(d)$ explains short- to medium-range self-excitation with a principled decay built into the prior malemshinitski2021 (i.e., without an explicit Gaussian envelope factor $e^{-\alpha d}$ multiplied the excitation function).
+The combination of (3) and (4) provides the necessary inductive bias for identifiability: repeated seasonal structure and trend are captured by $b(t)$, whereas $f(d)$ explains short- to medium-range self-excitation with a principled decay built into the prior Malem-Shinitski et al.
+(2022) (i.e., without an explicit Gaussian envelope factor $e^{-\alpha d}$ multiplied the excitation function).
 (a) $f(d)$ draws, $\beta=0.00$, $\ell_{f}=10$
@@
 Our inference procedure centers on the latent intensity function $l(t)$, defined in Equation (2), which we treat as the primary object of estimation.
-Since both the baseline $b(t)$ and the excitation kernel $f(d)$ are endowed with Gaussian process priors, the resulting $l(t)$ inherits a well-defined Gaussian process prior induced by these components (rasmussen2006gaussian, ).
+Since both the baseline $b(t)$ and the excitation kernel $f(d)$ are endowed with Gaussian process priors, the resulting $l(t)$ inherits a well-defined Gaussian process prior induced by these components (Rasmussen and Williams, 2006).
 Rather than estimating $b(t)$ and $f(d)$ jointly, we integrate them out analytically to form a collapsed prior over $l(t)$.
@@
 This formulation avoids the need to represent $b(t)$ and $f(d)$ explicitly during inference.
-This is analogous to GP-modulated Poisson processes in continuous time malemshinitski2021 , adapted here to the discrete-time Hawkes setting with a collapsed prior.
-Unlike malemshinitski2021 , which imposes an explicit exponential envelope together with a stationary RBF kernel for the excitation (i.e., contributions of the form $e^{-\alpha d}$ and covariance of the GP prior of the excitation function $k_{f}(d,d^{\prime})=\sigma_{f}^{2}\exp\!\{-(d-d^{\prime})^{2}/(2\ell_{f}^{2})\}$), we remove the deterministic envelope from the mean structure and encode attenuation and smoothness inside the GP prior for $f$.
+This is analogous to GP-modulated Poisson processes in continuous time Malem-Shinitski et al.
+(2022), adapted here to the discrete-time Hawkes setting with a collapsed prior.
+Unlike Malem-Shinitski et al.
+(2022), which imposes an explicit exponential envelope together with a stationary RBF kernel for the excitation (i.e., contributions of the form $e^{-\alpha d}$ and covariance of the GP prior of the excitation function $k_{f}(d,d^{\prime})=\sigma_{f}^{2}\exp\!\{-(d-d^{\prime})^{2}/(2\ell_{f}^{2})\}$), we remove the deterministic envelope from the mean structure and encode attenuation and smoothness inside the GP prior for $f$.
 3.2 MAP Objective
@@
     where $K_{U}$ is the stationary RBF kernel on the uniform grid.
-Because the grid is uniform, $K_{U}$ is (block-)Toeplitz/circulant and supports $\mathcal{O}(M\log M)$ MVMs via FFT (see xia2007toeplitz ).
+Because the grid is uniform, $K_{U}$ is (block-)Toeplitz/circulant and supports $\mathcal{O}(M\log M)$ MVMs via FFT (see Xia (2007)).
 Since $W$ is very sparse, $W\boldsymbol{v}$ and $W^{\top}\boldsymbol{v}$ cost $\mathcal{O}(T)$.
@@
 We evaluate GP-DHP and four benchmark parametric discrete-time Hawkes models on two real-world count datasets.
-For the benchmarks, we fix the excitation function to a negative-binomial form (a commonly used parametric form of the excitation function (white2012terrorism, )), whereas GP-DHP models the excitation nonparametrically via a Gaussian process prior.
+For the benchmarks, we fix the excitation function to a negative-binomial form (a commonly used parametric form of the excitation function (Porter and White, 2012)), whereas GP-DHP models the excitation nonparametrically via a Gaussian process prior.
 All benchmark models share the same negative-binomial excitation kernel:
@@
 $\Phi(d)=(1-p)^{d}p,$
-which has been used in recent infectious disease applications, including modeling COVID-19 mortality browning2021 .
-The more flexible negative binomial kernel has also been used in terrorism modeling white2012terrorism .
+which has been used in recent infectious disease applications, including modeling COVID-19 mortality Browning et al.
+(2021).
+The more flexible negative binomial kernel has also been used in terrorism modeling Porter and White (2012).
 Each model is evaluated using a one-step-ahead predictive log-likelihood (pLL) on a held-out test set.
@@
 4.2.1 U.S. Terrorism
-We analyze a 21-year time series of daily terrorist incidents within the United States, drawn from the Global Terrorism Database (GTD) gtd .
+We analyze a 21-year time series of daily terrorist incidents within the United States, drawn from the Global Terrorism Database (GTD) for the Study of Terrorism and to Terrorism) .
 The data spans from 1972 onward.
@@
 Self-exciting models offer a principled alternative.
-Prior work by (white2012terrorism, ) applied a discrete Hawkes process with a negative binomial excitation kernel to terrorism data in Southeast Asia, demonstrating how such models can recover meaningful structure in burst-prone time series and yield interpretable quantities such as volatility and resilience.
+Prior work by (Porter and White, 2012) applied a discrete Hawkes process with a negative binomial excitation kernel to terrorism data in Southeast Asia, demonstrating how such models can recover meaningful structure in burst-prone time series and yield interpretable quantities such as volatility and resilience.
 In our application, GP-DHP achieves the highest predictive log-likelihood on the test set, outperforming all parametric baselines (Table 2).
@@
 4.2.2 Cryptosporidiosis
-We analyze a 365-week series of weekly Cryptosporidiosis case counts, available via the tscount package in R tscount .
+We analyze a 365-week series of weekly Cryptosporidiosis case counts, available via the tscount package in R Liboschik et al.
+(2017).
 Cryptosporidiosis is a gastrointestinal disease caused by the protozoan parasite Cryptosporidium, which is typically spread through ingestion of contaminated water.
 While low-level background transmission persists in many regions, the disease is most notable for its sporadic and localized outbreaks.
-One of the most severe documented events occurred in Milwaukee in 1993, where over 400,000 residents were affected due to a contaminated municipal water supply hunter2003epidemiology ; yoder2010cryptosporidiosis .
+One of the most severe documented events occurred in Milwaukee in 1993, where over 400,000 residents were affected due to a contaminated municipal water supply Hunter et al.
+(2003); Yoder et al.
+(2010).
 These outbreak dynamics are characterized by sharp, short-term increases in reported cases, followed by a rapid return to low endemic levels.
-Such temporal patterns are challenging for classical autoregressive count models, such as Poisson or INGARCH-type models fokianos2009poisson ; fokianos2011log , which assume dependence based on past counts but do not directly model the self-exciting nature of outbreaks.
+Such temporal patterns are challenging for classical autoregressive count models, such as Poisson or INGARCH-type models Fokianos et al.
+(2009); Fokianos and Tjøstheim (2011), which assume dependence based on past counts but do not directly model the self-exciting nature of outbreaks.
 These models often struggle to distinguish between sustained background intensity and transient, event-driven clustering.
@@
 This separation provides epidemiologically meaningful insights: the baseline reflects ongoing exposure or environmental risk, while the excitation term captures outbreak-driven propagation.
-Recent work by (browning2021, ) demonstrated the use of discrete Hawkes models with geometric excitation kernels for modeling daily COVID-19 mortality in multiple countries.
+Recent work by (Browning et al., 2021) demonstrated the use of discrete Hawkes models with geometric excitation kernels for modeling daily COVID-19 mortality in multiple countries.
 The geometric kernel is a special case of the negative binomial excitation used here (specifically, when $r=1$), and has proven effective for modeling short-memory self-excitation in epidemic settings.